# Chain of Preference Optimization: Improving Chain-of-Thought Reasoning in LLMs

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) have shown that constructing reasoning chains is critical to improving their problem-solving capabilities [40, 39, 54, 47, 53]. A representative method is chain-of-thought (CoT) [40], which prompts LLMs to generate intermediate reasoning steps, i.e., thoughts, thereby constructing explicit reasoning paths (as depicted in Figure 1(a)). While straightforward and intuitive, recent research observes that CoT can often overlook optimal reasoning paths and exhibit an unconscious style of answering due to its single-path focus [48, 7]. To foster a more deliberate and conscious reasoning style, Yao et al. [48] propose tree-of-thought (ToT), which generates multiple branching thoughts at each step of the reasoning process and conducts self-evaluation for pruning and planning to search for an optimal reasoning path (as shown in Figure 1(b)). However, despite improving reasoning quality, ToT significantly increases computational complexity, which limits its practical application. This raises the question: Can the strategic depth of ToT be integrated into CoT to enhance its effectiveness while maintaining efficiency?
@@
In this paper, we identify a limitation in these approaches: they overlook the non-optimal reasoning thoughts generated during the tree-search process, which naturally provides additional preference information. Specifically, ToT inherently generates multiple alternative thoughts at each reasoning step, and pruning is performed according to their evaluated qualities. Intuitively, this tree-search process constitutes a preference over all intermediate thought candidates—thoughts appearing in the best-discovered reasoning path should be preferred over those that do not. Moreover, this could shed even more insights than the final best-discovered reasoning path, as non-optimal reasoning paths (and thus preferences) exist at each step in the tree-search.
@@ 2 Related Work
LLMs have been shown to perform better when prompted to engage in multi-step reasoning [40, 39, 54]. Many studies have focused on improving the generated reasoning paths by post-editing [53] or accessing external knowledge [54, 31]. A distinct approach, more relevant to our interests, transforms the linear reasoning structure into a non-linear format such as a tree or graph [49, 27, 48, 7, 17], which combines thought evaluation with search algorithms like depth-first search (DFS) [12]. Different from our proposed CPO, these methods require searching during inference, which significantly increases latency.
@@
Reinforcement learning (RL) has increasingly been applied to LLMs by treating them as RL agents for alignment with human feedback [28, 43, 5, 42]. Recent advances demonstrate the potential of using LLMs for self-generating data to augment fine-tuning processes [6, 50, 11]. For instance, reinforced self-training methods [16, 1, 45] introduce mechanisms to curate new high-quality examples and iteratively enrich the training dataset for enhancing model performance. Nevertheless, these methods typically rely on either an external reward model [45, 1] or labeled data [16]. In contrast, approaches like self-rewarding [11, 24] utilize LLMs themselves to evaluate the generated content, aligning more closely with our method. However, these strategies still require initial seed data [11, 24], necessitating human annotation. Our work differs from previous methods as it does not rely on any ground-truth data, allowing LLMs to self-learn from their own feedback. Additionally, our approach constructs feedback in a chain fashion, focusing on reasoning steps, an aspect overlooked by prior works.
@@
Monte Carlo tree-search (MCTS) is a robust algorithm for navigating complex decision-making environments, commonly employed in strategic board games such as AlphaGo [8, 23, 41, 10, 34]. MCTS methodically constructs a search tree, balancing exploration and exploitation, simulates various outcomes, and updates utility estimates based on these simulations. Recent studies have shown that MCTS can enhance the decoding process in LLMs [14, 25, 13, 17, 36]. However, the primary challenge with MCTS for LLM is the high latency during inference. While some approaches have attempted to optimize LLMs by leveraging reasoning paths identified through MCTS [14, 36], these methods still rely on labeled data and require separate policy and value models to explore and evaluate potential moves at the tree’s leaves. In contrast, our CPO approach eliminates the need for human annotations and simplifies the tuning of LLMs without the necessity for additional models.
@@ 4 Our Method: Chain of Preference Optimization
Figure 2: The framework of our CPO method. The left part illustrates the process of generating, evaluating, and pruning thoughts, while the right part demonstrates the collection of preference thoughts. The shaded path represents the final selected reasoning path. Thoughts marked with a trophy indicate preferred data, and their sibling nodes are dispreferred.
Unlike previous methods that train LLMs to learn the complete reasoning path [21, 38, 14, 36], our approach leverages the preferences over thoughts generated at each reasoning step, which are often discarded in prior works. Our key insight is that non-optimal thoughts generated during the tree-search process in ToT provide valuable preference information that can enhance LLM’s reasoning ability. A major advantage of our method is that it utilizes this supervision only during training, thereby avoiding high inference latency. Our approach consists of two components: synthesizing the chain of preference thoughts (i.e., the preference thoughts in a chain fashion) and training with the CPO objective.
@@ 4.1 Synthesizing the Chain of Preference Thoughts
To minimize the effects of randomness and bias, we shuffle the order of demonstration examples [3] and repeatedly sample the generated justification and evaluation results. We then calculate the average score for the state $s^{(j)}_{i}$. The general guideline prompt for the evaluation is as follows: Evaluate whether the thought helps in partially or directly answering the original question (likely/impossible).
@@ 4.2 Training with the CPO Objective
Thus, the objective function for the chain of preference thoughts can be formulated as follows:
@@ 5 Experiments
In this section, we empirically validate that CPO improves the reasoning ability of the base models and uncover several insightful findings.
@@ 5.1 Settings
We focus our research on three types of reasoning tasks: Question Answering (QA), Fact Verification, and Arithmetic Reasoning. For QA, we conduct experiments on three widely used datasets: Bamboogle [31], WikiMultiHopQA [18], and HotpotQA [44]. For fact verification, we use three datasets: Fever [35], Feverous [2], and Vitaminc [33]. For arithmetic reasoning, we test on the SVAMP dataset [30]. We use 4-shot prompting for each dataset, with CoT demonstrations manually constructed by previous works [40, 31, 39]. Detailed experimental configurations can be found in Appendix A. For evaluation metrics, we report the accuracy and the average latency of generating the answer per instance.
@@ 5.2 Overall Results on Reasoning
CPO has lower latency than ToT while maintaining comparable performance.
@@
CPO surpasses TS-LLM on average.
Despite both CPO and TS-LLM using ToT to generate training data (where our implementation of ToT remains consistent), CPO exhibits an average improvement of $2.7\%$ and reaches a maximum increase of $10.3\%$. A key factor behind this performance is the CPO’s ability to fully utilize the ToT reasoning process. Specifically, CPO effectively leverages both selected and unselected thoughts at each reasoning step, whereas TS-LLM only uses information from the selected paths, offering CPO with a clear advantage. A detailed discussion of the effectiveness of CPO is presented in Section 5.3.
@@ 5.3 Component-wise Evaluations
Effect of the number of training data.
@@
Sensitivity to data mixture.
@@ 6 Analysis
We explore the impact of dispreferred thoughts on model performance by gradually incorporating these thoughts into the training data. Initially, we introduce dispreferred thoughts for their corresponding preferred counterparts and apply CPO to this segment of the data. For preferred thoughts without dispreferred counterparts, we implement SFT on these data. Consequently, the percentage of dispreferred thoughts incorporated can also be viewed as the proportion of data processed using CPO. We adjust the inclusion percentage of dispreferred thoughts from $0\%$ to $100\%$. An inclusion of $0\%$ indicates that we utilize SFT solely on the preferred thoughts, i.e., the baseline TS-SFT. Conversely, an inclusion of $100\%$ signifies our CPO, where the entire dataset includes paired preferred and dispreferred thoughts. As shown in Figure 3(d), we find that increasing the percentage of dispreferred data inclusion consistently improves model performance. This suggests that dispreferred thoughts are beneficial during the optimization process, highlighting the importance of leveraging both preferred and dispreferred thoughts for enhancing the model’s reasoning capabilities.
@@ 7 Conclusion
For future work, we aim to combine CPO with other reasoning algorithms, such as graph-of-thoughts [7]. Additionally, we are interested in exploring the potential of using a weak LLM to evaluate a strong one within the CPO framework, aiming to achieve weak-to-strong alignment [9].
@@ Appendix A Detailed Experiment Configurations
To maintain a reasonable budget, especially given the high computational demand of ToT, we limit each dataset to a maximum of 300 test samples through random sampling. For datasets that contain less than 300 test samples, we instead use all available samples. For training, we randomly select less than 300 instances from each dataset to construct the preference data pairs, without using the ground-truth labels. This is because we observe that more number of training data does not lead to performance improvement as shown in Section 5.3.
Appendix B CPO Benefits from Iterative Learning.
Inspired by the iterative improvements achieved in previous research [50, 51], in this section, we explore whether CPO can be further improved by iterative learning. Specifically, we try two distinct iterative training strategies: 1) SFT+CPO: in iter=0, Start with a base LLM that has not been fine-tuned at all; in iter=1, SFT the base LLM on the reasoning path selected by ToT (base model); in subsequent iterations (iteration >1), Continue to fine-tune the model using the CPO method, based on the chain of preference thoughts constructed by the model in the previous iterations. and 2) CPO only: in iter=0, same as iter=0 in SFT+CPO; in subsequent iterations (iteration >0): Only use the CPO method for training in all iterations, similar to the approach in SFT+CPO after the first iteration.

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) have shown that constructing reasoning chains is critical to improving their problem-solving capabilities [1, 2, 3, 4, 5, 6, 7]. A representative method is chain-of-thought (CoT) [1], which prompts LLMs to generate intermediate reasoning steps, i.e., thoughts, thereby constructing explicit reasoning paths (as depicted in Figure 1(a)). While straightforward and intuitive, recent research observes that CoT can often overlook optimal reasoning paths and exhibit an unconscious style of answering due to its single-path focus [8, 9]. To foster a more deliberate and conscious reasoning style, Yao et al. [8] propose tree-of-thought (ToT), which generates multiple branching thoughts at each step of the reasoning process and conducts self-evaluation for pruning and planning to search for reasoning paths (as shown in Figure 1(b)). However, despite improving reasoning quality, ToT significantly increases computational complexity, which limits its practical application. This raises the question: Can the strategic depth of ToT be integrated into CoT to enhance its effectiveness while maintaining efficiency?
@@
In this paper, we identify a limitation in these approaches: they overlook the non-optimal reasoning thoughts generated during the tree-search process, which naturally provides additional preference information. Specifically, ToT inherently generates multiple alternative thoughts at each reasoning step, and pruning is performed according to their evaluated qualities. This tree-search process constitutes a preference over all intermediate thought candidates—thoughts appearing in the best-discovered reasoning path are preferred over those that do not. Moreover, this could shed even more insights than the final best-discovered reasoning path, as non-optimal reasoning paths (and thus preferences) exist at each step in the tree-search.
@@ 2 Related Work
LLMs have been shown to perform better when prompted to engage in multi-step reasoning [1, 2, 3]. Many studies have focused on improving the generated reasoning paths by post-editing [16] or accessing external knowledge [3, 17]. A distinct approach, more relevant to our interests, transforms the linear reasoning structure into a non-linear format such as a tree or graph [18, 19, 20, 8, 9, 21], which combines thought evaluation with search algorithms like depth-first search (DFS) [22]. Different from our proposed CPO, these methods require searching during inference, which significantly increases latency.
@@
Reinforcement learning (RL) has increasingly been applied to LLMs by treating them as RL agents for alignment with human feedback [23, 24, 25, 26]. Recent advances demonstrate the potential of using LLMs for self-generating data to augment fine-tuning processes [27, 28, 29, 30, 31, 32]. For instance, reinforced self-training methods [33, 34, 35, 36, 37] introduce mechanisms to curate new high-quality examples and iteratively enrich the training dataset for enhancing model performance. Nevertheless, these methods typically rely on either an external reward model [35, 34] or labeled data [33]. In contrast, approaches like self-rewarding [29, 38] utilize LLMs themselves to evaluate the generated content, aligning more closely with our method. However, these strategies still require initial seed data [29, 38], necessitating human annotation. Our work differs from previous methods as it does not rely on any ground-truth data, allowing LLMs to self-learn from their own feedback. Additionally, our approach constructs feedback in a chain fashion, focusing on reasoning steps, an aspect overlooked by prior works.
@@
Monte Carlo tree-search (MCTS) is a robust algorithm for navigating complex decision-making environments, commonly employed in strategic board games such as AlphaGo [39, 40, 41, 42, 43]. MCTS methodically constructs a search tree, balancing exploration and exploitation, simulates various outcomes, and updates utility estimates based on these simulations. Recent studies have shown that MCTS can enhance the decoding process in LLMs [11, 44, 45, 21, 12]. However, the primary challenge with MCTS is the high latency during inference, particularly in difficult reasoning tasks [46, 47]. While some approaches have attempted to optimize LLMs by leveraging reasoning paths identified through MCTS [11, 12], these methods still rely on labeled data and require separate policy and value models to explore and evaluate potential moves at the tree’s leaves. In contrast, our CPO approach eliminates the need for human annotations and simplifies the tuning of LLMs without the necessity for additional models.
@@ 4 Our Method: Chain of Preference Optimization
Figure 2: The framework of our CPO method. The left part illustrates the process of generating, evaluating, and pruning thoughts, while the right part demonstrates the collection of preference thoughts. The shaded path represents the final selected reasoning path. Thoughts marked with a trophy indicate preferred data, while sibling nodes without a trophy are marked as dispreferred.
Unlike previous methods that train LLMs to learn the complete reasoning paths [10, 50, 11, 12], our approach leverages the preferences over thoughts generated at each reasoning step, which are often discarded in prior works. Our key insight is that non-optimal thoughts generated during the tree-search process in ToT provide valuable preference information that can enhance LLM’s reasoning ability. A major advantage of our method is that it utilizes this supervision only during training, thereby avoiding high inference latency. Our approach consists of two components: synthesizing the chain of preference thoughts (i.e., the preference thoughts in a chain fashion) and training with the CPO objective.
@@ 4.1 Synthesizing the Chain of Preference Thoughts
The prompt template used in our evaluation consists of two parts: (1) the general guidelines, and (2) task-specific demonstration examples. To minimize the effects of randomness and bias, we shuffle the order of demonstration examples [51] and repeatedly sample the generated justification and evaluation results. We then calculate the average score for the state $s^{(j)}_{i}$. The general guideline prompt for the evaluation is as follows: Evaluate whether the thought helps in partially or directly answering the original question (likely/impossible).
@@ 4.2 Training with the CPO Objective
Thus, the objective function for CPO can be formulated as follows:
@@ 5 Experiments
In this section, we empirically validate that CPO improves the reasoning ability of the base model, and uncover several insightful findings.
@@ 5.1 Settings
We focus our research on three types of reasoning tasks: Question Answering (QA), Fact Verification, and Arithmetic Reasoning. For QA, we conduct experiments on three widely used datasets: Bamboogle [17], WikiMultiHopQA [52], and HotpotQA [53]. For fact verification, we use three datasets: Fever [54], Feverous [55], and Vitaminc [56]. For arithmetic reasoning, we test on the SVAMP dataset [57]. Our choice of tasks was driven by the performance of the models using the ToT method, which showed improvements in QA, fact verification, and arithmetic reasoning. We use 4-shot prompting for each dataset, with CoT demonstrations manually constructed by previous works [1, 17, 2]. Detailed experimental configurations can be found in Appendix A. For evaluation metrics, we report the accuracy and the average latency of generating the answer per instance. More metrics can be found in Appendix B.
@@ 5.2 Overall Results on Reasoning
CPO has a lower latency than ToT but comparable performance.
@@
CPO surpasses TS-SFT on average.
Despite both CPO and TS-SFT using ToT to generate training data (where our implementation of ToT remains consistent), CPO exhibits an average improvement of $2.7\%$ and reaches a maximum increase of $10.3\%$. A key factor behind this performance is the CPO’s ability to fully utilize ToT’s reasoning process. Specifically, CPO effectively leverages both selected and unselected thoughts at each reasoning step, whereas TS-SFT only uses information from the selected paths, offering CPO with a clear advantage. A detailed discussion of the effectiveness of CPO is presented in Section 5.3.
@@ 5.3 Component-wise Evaluations
Effect of numbers of training data.
@@
Sensitivity of data mixture.
@@ 6 Analysis
We explore the impact of dispreferred thoughts on model performance by gradually incorporating these thoughts into the training data. Initially, we introduce dispreferred thoughts for their corresponding preferred counterparts and apply CPO to this segment of the data. For preferred thoughts without dispreferred counterparts, we implement SFT on these data. Consequently, the percentage of dispreferred thoughts incorporated can also be viewed as the proportion of data processed using CPO. We adjust the inclusion percentage of dispreferred thoughts from $0\%$ to $100\%$. An inclusion of $0\%$ indicates that we utilize SFT solely on the preferred thoughts, i.e., the baseline TS-SFT. Conversely, an inclusion of $100\%$ signifies our CPO, where the entire dataset includes paired preferred and dispreferred thoughts.
@@
As shown in Figure 3(d), we find that increasing the percentage of dispreferred data inclusion consistently improves model performance. This suggests that dispreferred thoughts are beneficial during the optimization process, highlighting the importance of leveraging both preferred and dispreferred thoughts for enhancing the model’s reasoning capabilities.
@@ 7 Conclusion
In future work, we plan to integrate CPO with other reasoning algorithms, such as Graph-of-Thoughts [9] and AlphaZero-like tree search [11]. Furthermore, we intend to explore the potential of using a weaker LLM to evaluate a stronger one within the CPO framework, facilitating weak-to-strong alignment [64].
Acknowledgement
We thank Cunxiao Du for the insightful derivation of the Longest Common Prefix (LCP) gradient cancellation and for the valuable contributions to the discussions throughout this work. We also appreciate Fangkai Jiao’s helpful feedback. Additionally, we thank the anonymous reviewers for their constructive comments and suggestions that helped improve this paper.
@@ Appendix A Detailed Experiment Configurations
To maintain a reasonable budget, especially given the high computational demand of ToT, we limit each dataset to a maximum of 300 test samples through random sampling. For datasets that contain less than 300 test samples, we instead use all available samples. For training, we randomly select less than 300 instances from each dataset to construct the preference data pairs, without using the ground-truth labels. This is because we observe that more number of training data does not lead to performance improvement as shown in Section 5.3. In addition, in our experiments, approximately 200 samples (e.g., questions in the QA task) on average could generate about 6,531 preference pairs, suggesting that our CPO requires only a small number of samples by design. Constructing preference data is a time-intensive process. The choice of 300 samples represents a practical trade-off between efficiency and effectiveness, allowing us to manage resources effectively while still achieving noticeable improvements.
Appendix B Additional Experiments
CPO benefits from iterative learning.
Inspired by the iterative improvements achieved in previous research [28, 78], in this section, we explore whether CPO can be further improved by iterative learning. Specifically, we try two distinct iterative training strategies: 1. SFT+CPO: in iter=0, Start with a base LLM that has not been fine-tuned at all; in iter=1, SFT the base LLM on the reasoning path selected by ToT (base model); in subsequent iterations (iteration >1), Continue to fine-tune the model using the CPO method, based on the chain of preference thoughts constructed by the model in the previous iterations. and 2. CPO only: in iter=0, same as iter=0 in SFT+CPO; in subsequent iterations (iteration >0): Only use the CPO method for training in all iterations, similar to the approach in SFT+CPO after the first iteration.
@@ Appendix B Additional Experiments
Comparasion with ReST and self-rewarding baseline.
The settings of ReST [33] and self-rewarding [28] are different from ours, as discussed in Section 2. These methods rely on either external reward models (ReST) or labeled data (self-rewarding), which makes them not directly comparable to our approach. To ensure a fair comparison with our CPO method, we prompted the LLM itself in the same way as our CPO to serve as the reward model for ReST and as the labeled data annotator for self-rewarding, respectively. As shown in Table 4, the results indicate that, on average, our CPO method surpasses both ReST and self-rewarding under this fair comparison setting.
Model Bam. 2wiki Hotpotqa Fever Feverous Vitaminc Svamp Average
LLaMA2-7b
Rest 30.4 24.0 22.3 45.5 43.9 51.7 42.3 37.2
SelfR 31.2 25.3 21.0 48.8 44.7 51.3 43.0 37.5
CPO 32.0 29.7 24.0 53.2 49.0 52.7 46.0 40.9
LLaMA2-13b
Rest 48.0 28.3 28.7 46.8 48.0 50.2 44.3 42.0
SelfR 48.0 29.0 30.0 47.8 48.5 51.0 45.3 42.8
CPO 52.0 30.3 30.3 49.2 50.7 54.0 50.0 45.2
Mistral-7b
Rest 43.2 26.7 27.4 59.5 48.3 49.7 63.3 45.4
SelfR 44.0 28.0 28.1 58.2 48.0 50.0 65.0 45.9
CPO 45.6 31.7 29.4 59.9 54.0 53.7 69.3 49.1
Table 4: Performance comparison across different datasets.
More metrics on QA dataset.
We include F1 scores for the three QA tasks in Table 5. The results show that the F1 performance aligns well with the corresponding accuracy for each task.
LLaMA2-7b LLaMA2-13b Mistral-7b
ToT CoT SFT CPO ToT CoT SFT CPO ToT CoT SFT CPO
Bam. 29.5 28.1 29.3 31.8 43.0 40.9 41.4 43.0 48.1 42.2 41.9 47.5
2wiki. 26.5 23.6 25.0 26.0 30.7 26.8 27.2 28.3 25.5 24.3 27.3 27.0
hot. 23.4 20.5 22.9 24.7 33.3 31.3 31.3 32.0 29.9 25.9 26.8 28.0
Table 5: F1 scores on QA datasets.
Ablation studies across datasets and models.
Figure 67, and 8 provide ablations and analysis across various models and datasets. The observed trends remain generally consistent across these different settings.
Figure 6: Effect of the number of instances in generating paired thoughts.
Figure 7: Effect of dispreferred thoughts in optimization.
(a) Effect of dispreferred thoughts selection.
(b) Effect of per-step preference supervision.
Figure 8: Component-wise Evaluations and Analysis.
Illustrative examples of the reasoning paths preferred by CPO.
Table 6 presents examples demonstrating that the paths favored by CPO align more closely with those selected by ToT than by CoT, indicating a higher reasoning quality in CPO’s chosen paths.
Question 1: When did the author of Annabel Lee enlist in the army?
CoT: Step 1, when did Edgar Allan Poe enlist in the army? Edgar Allan Poe enlisted in the army on May 27, 1827. Step 2, when did Edgar Allan Poe die? Edgar Allan Poe died on October 7, 1849. Step 3, so the final answer is: October 7, 1849.
ToT: step 1, when did edgar allan poe enlist in the army? edgar allan poe enlisted in the army on august 16, 1827.step 2, what year did edgar allan poe enlist in the army? edgar allan poe enlisted in the army in 1827.step 3, so the final answer is: 1827.
CPO: step 1, when did Edgar Allan Poe die? Edgar Allan Poe died on October 7, 1849. Step 2, when did Edgar Allan Poe enlist in the army? Edgar Allan Poe enlisted in the army on May 27, 1827. Step 3, so the final answer is: May 27, 1827.
Question 2: What is the nickname of the easternmost U.S. state?
CoT: Step 1, who is the scientist at MIT that won the Queen Elizabeth Prize for Engineering in 2013? The scientist at MIT that won the Queen Elizabeth Prize for Engineering in 2013 was Daniela Rus. Step 2, who is the father of Daniela Rus? Daniela Rus’s father is Vladimir Rus. Step 3, so the final answer is: Vladimir Rus.
ToT: step 1, what is the name of the easternmost u.s. state? maine is the easternmost u.s. state.step 2, what is the nickname of maine? maine’s nickname is ’the pine tree state.
CPO: Step 1, what is the name of the easternmost U.S. state? The easternmost U.S. state is Maine. Step 2, what is the nickname of Maine? The nickname of Maine is The Pine Tree State. Step 3, so the final answer is: The Pine Tree State.
Table 6: Illustrative examples of the reasoning paths preferred by CPO. Similar paths between ToT and CPO are in yellow.